The driver of a speeding empty truck slams on the brakes and slows down to a stop through a distance d. On a second trial, the truck carries a load that doubles its mass. What distance will the loaded truck cover to come to rest, when it is moving at the same speed and the brakes are slammed?
1
Expert's answer
2020-10-15T03:13:54-0400

Empty truck:

"d=\frac{v^2}{2a}"

"F=ma=m\frac{v^2}{2d}"


The truck with load ("m'=3m)" :

"d'=\frac{v'^2}{2a'}"

"F'=3ma'=3m\frac{v^2}{2d'}"


Then "F=F'"


"m\frac{v^2}{2d}=3m\frac{v^2}{2d'}"


"\frac{3}{d'}=\frac{1}{d}" where "d'=3d"